# Entropy Meets Importance: A Unified Head Importance–Entropy Score for Stable and Efficient Transformer Pruning

### v1

@@ 1 Introduction
However, HIS-based methods often exhibit limited stability in their performance. For clarity, prior works (Bair et al., 2024; Blanchet et al., 2024) motivate treating stability as a practical surrogate for robustness—namely, a model’s resilience to input perturbations and pruning-induced distributional shifts. Such stability is crucial in real-world deployments where distribution shifts are common and aggressive compression is often required. In our observations, this instability appears to stem from two key factors. First, existing HIS-based methods solely rely on the loss gradient with respect to each head’s output, which fails to capture token-level attention allocation or its alignment with the task’s empirical distribution. Consequently, a concentrated head and a diffuse head can yield similar important scores, concealing their functionally distinct roles on the task-specific data manifold. Second, a uniform, layer-agnostic criterion precludes layer-specific adaptation despite evidence that different layers require distinct attention behaviors (Artzy and Schwartz, 2024). Lacking such layer-specific characteristics often results in imbalanced pruning—preserving redundant heads in some layers while removing functionally critical ones in others. This imbalance not only degrades accuracy but also undermines stability, leading to unpredictable performance fluctuations across inputs or compression levels, particularly under aggressive pruning ratios.
@@ 2 Background
Attention head pruning. To compress large language models efficiently, structured pruning methods (Han et al., 2015; Wang et al., 2019; Hou et al., 2020a; Ma et al., 2023; Ashkboos et al., 2024), which remove specific architectural components from Transformer models, have been widely adopted. Among these, attention head pruning has gained traction. This is largely because it directly reduces attention FLOPs and KV-cache memory while preserving the layer topology, thereby simplifying checkpoint compatibility and serving integration. Consequently, large‐scale studies adopt head-level pruning as a practical axis in LLM compression pipelines (Jaradat and et al., 2024; Muralidharan et al., 2024) [^1] For more detailed discussions on related work, please refer to Appendix A.. Attention head pruning removes selected heads from a trained Transformer’s multi‑head attention with minimal impact on end‑task performance (Vaswani et al., 2017). A widely adopted criterion is the HIS of Michel et al. (2019), which introduces mask variables $m_{h}\in\{0,1\}$ multiplying the output of head $h$ and defines importance as the expected first-order loss increase under masking:
@@ 3 Motivation
Pruning Transformer models is most commonly driven by gradient-based criteria, such as HIS and variants used in recent pruning frameworks (e.g., LLM-Pruner) [Michel et al., 2019, Ma et al., 2023]. While gradient-based methods are often effective at moderate sparsity, they exhibit sharp accuracy degradation once the pruning ratio exceeds a certain threshold, as shown in Fig. 1 (a). Such sharp-drops have been widely observed across various attention variants and tasks [Ma et al., 2023, Mao et al., 2023, Ghattas et al., 2025], underscoring the generality of this phenomenon.
We focus on this “sharp-drop” regime and contrast two groups of heads. The first group consists of low-HIS heads that are pruned during the sharp-drop of accuracy. The second group consists of high-HIS heads that remain unpruned. The attention score heatmap in Fig. 1 (b) reveals that some pruned heads (red table) assign high attention scores—i.e., the weights computed by the softmax over token-token similarity that indicate how strongly a token attends to another—to sentiment-discriminative tokens (tokens relevant for classification). Nonetheless, these heads are pruned due to their low HIS values and end up causing the sharp accuracy drop observed in Fig. 1 (a). In contrast, some unpruned heads (gray table) often allocate strong attention to non-informative tokens. These heads, however, have high HIS values and thus remain unpruned, though they contribute little to overall model quality. These analysis results demonstrate that the gradient-based $\mathrm{HIS}$ is insufficient to capture the token-level attention score distributions (the detailed mathematical analysis is provided in Section 4.2.1), thereby resulting in suboptimal pruning decisions for heads focusing on decisive tokens.
The sharp-drop observed in pruning can be interpreted as a collapse of structural diversity in attention behaviors, caused by the elimination of heads that concentrate on decisive tokens. To capture and prevent such collapse of structural diversity in attention, we employ attention entropy, a measure widely used to prevent policy collapse in reinforcement learning [Bharadhwaj et al., 2020, Liu et al., 2021a, Wang et al., 2025]. As shown in the bottom-right plots of Fig. 1 (b), incorporating AE helps retain low-entropy heads by recognizing their concentrated focus on decisive tokens, thereby preventing them from being pruned and mitigating the sharp-drop in accuracy.
@@ 4 Proposed Method
Fig. 2 provides an overview of our proposed Head Importance–Entropy Score (HIES). Fig. 2 (a) outlines the Transformer architecture, where HIS computes the importance score of each of the $h$ attention heads within a Transformer block. Fig. 2 (b) displays the heatmap of the first-order, loss-based HIS. Combining the normalized HIS and AE yields the HIES heatmap in Fig. 2 (c), which integrates complementary signals provide a more stable assessment of attention heads across layers. This design captures the key intuition of HIES: HIS can be reinforced by AE, providing a more robust basis for pruning head selections. Section 4.1 formalizes HIES, and Section 4.2 develops a risk-decomposition analysis that clarifies the complementary roles of HIS and AE and motivates HIES as a robust criterion for head selection in pruning.
@@ 4.1 Head Importance Entropy Score
Min-Max Normalization Directly comparing raw HIS and AE is inherently problematic, as the two metrics reside on different scales and encode distinct types of signals. To enable meaningful integration and ranking, we apply min–max normalization to both metrics, rescaling their values to the interval $[0,1]$: $\widehat{\text{HIS}}_{h}=\frac{\text{HIS}_{h}-\min(\text{HIS})}{\max(\text{HIS})-\min(\text{HIS})},\quad\widehat{\text{AE}}_{h}=\frac{\text{AE}_{h}-\min(\text{AE})}{\max(\text{AE})-\min(\text{AE})}.$ This distribution-agnostic normalization improves cross-criterion interpretability; lower normalized scores denote higher pruning priority. Prior studies show min–max scaling outperforms $z$-score standardization in stability and reproducibility across diverse tasks [de Amorim et al., 2022, Lima and Souza, 2023].
@@ 4.2.1 Loss-Increase Control via Head Importance (HIS)
$\mathrm{HIS}_{h}~:=~\mathbb{E}_{x\sim\mathcal{D}}\!\left|\frac{\partial\mathcal{L}(x)}{\partial m_{h}}\right|~=~\mathbb{E}_{x\sim\mathcal{D}}\!\left|\,\big\langle\nabla_{A_{h}(x)}\mathcal{L}(x),\,A_{h}(x)\big\rangle_{F}\,\right|~=~\mathbb{E}_{\mathcal{D}}\!\Bigl[\,\left|\,\langle\nabla_{A_{h}}\mathcal{L},\;A_{h}\rangle_{F}\,\right|\,\Bigr].$ (1)
@@
Lemma 1 (Loss-increase upper bound under head masking).
Let $\beta_{y}:=\|\nabla_{\mathbf{y}}^{2}\mathcal{L}\|_{2}$. For any mask variables $\{m_{h}\}_{h=1}^{|\mathrm{H}|}$, $\Delta\mathcal{L}~:=~\mathbb{E}_{\mathcal{D}}\!\bigl[\mathcal{L}(\mathbf{y}+\delta\mathbf{y})-\mathcal{L}(\mathbf{y})\bigr]~\leq~\sum_{h=1}^{|\mathrm{H}|}(1-m_{h})\,\mathrm{HIS}_{h}~+~\frac{\beta_{y}}{2}\sum_{h=1}^{|\mathrm{H}|}(1-m_{h})\,\|A_{h}\|_{F}^{2}.$ Moreover, under binary (sigmoid) cross-entropy we have $\|\nabla_{\mathbf{z}}^{2}\mathcal{L}\|_{2}\leq\tfrac{1}{4}$; with the linear projection $\mathbf{z}=\mathbf{y}W^{O}$, this yields $\beta_{y}\leq\tfrac{1}{4}\,\|W^{O}\|_{2}^{2}$, hence
@@
Eq. equation 2 shows that, for a fixed pruning fraction $\rho=\tfrac{1}{\mathrm{|H|}}\sum_{h}(1-m_{h})$, selecting heads with the smallest $\mathrm{HIS}_{h}$ minimizes the dominant first-order term, while the quadratic term is controlled by $\|W^{O}\|_{2}$ (or blockwise norms) and token-averaged activations. Under standard normalization, the quadratic contribution is typically dominated by the first-order term (cf. Appendix B.1.5), justifying the use of $\mathrm{HIS}_{h}$ as a practical surrogate importance for head pruning.
However, since $A_{h}=\mathrm{Attn}_{h}V_{h}$, $\mathrm{HIS}_{h}$ in Eq. equation 1 depends solely on $A_{h}$ and $\nabla_{A_{h}}\mathcal{L}$, without explicitly incorporating the distribution $\mathrm{Attn}_{h}$. Consequently, two heads with very different attention patterns (e.g., sharply focusing on tokens versus spreading over tokens) can yield similar $\mathrm{HIS}_{h}$ whenever the resulting $A_{h}$ is comparable. This is supported by our empirical analysis in Section 3.
@@ 4.2.2 Generalization Gap and Attention Entropy (AE)
$\mathrm{AD}_{h}(x)$ $=\frac{1}{n(x)\,\log n(x)}\sum_{t=1}^{n(x)}\Bigl(\log n(x)-H\!\bigl(\bm{\alpha}^{(h)}_{t}(x)\bigr)\Bigr)\;=1-\mathrm{AE}_{h}(x)\in[0,1].$ (3)
@@ 4.2.3 Risk Upper Bound and HIES Minimization
$\mathcal{R}(m)\;:=\;\sum_{h=1}^{|\text{H}|}(1-m_{h})\,\text{HIES}_{h}\vskip-4.2679pt$ (5)
@@
$\min_{m\in\{0,1\}^{|\text{H}|}}\;\sum_{h=1}^{|\text{H}|}(1-m_{h})\,\text{HIES}_{h}\quad\text{s.t.}\;\sum_{h=1}^{|\text{H}|}m_{h}\;=\;k.\vskip-2.84526pt$ (6)
Lemma 2 (Optimality).
Selecting the $k$ heads with the largest HIES values (equivalently, pruning the $|\text{H}|-k$ heads with the smallest HIES values) yields the globally optimal mask $m^{*}$ that minimizes equation 5 subject to equation 6.
@@ 4.2.4 Orthogonality and Complementarity
$u_{h}\;:=\;\operatorname{sign}\!\bigl(\bm{\alpha}^{(h)\!\top}g_{h}\bigr)\,g_{h},\qquad v_{h}\;:=\;\mathbf{1}+\log\bm{\alpha}^{(h)},\qquad\tilde{u}_{h}:=P\,u_{h},\;\;\tilde{v}_{h}:=P\,v_{h},$
where $P:=I-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\!\top}$ projects onto $\{w:\mathbf{1}^{\!\top}w=0\}$. Assume $\mathrm{Cov}(\tilde{u}_{h},\tilde{v}_{h})=0$ (the cross-covariance matrix is zero). If, in addition, either $\mathbb{E}_{\mathcal{S}}[\tilde{u}_{h}]=0$ or more generally $\langle\mathbb{E}_{\mathcal{S}}[\tilde{u}_{h}],\,\mathbb{E}_{\mathcal{S}}[\tilde{v}_{h}]\rangle=0$, then
@@
i.e., the two gradient directions are orthogonal in expectation. [^3] Detailed derivations and preliminaries are deferred to Appendix B.4.
@@ 5.1 Experimental Setup
Model. We use publicly available $\text{BERT}_{\text{base}}$ checkpoints that have been fine-tuned and released by prior work [Devlin et al., 2019], and $\text{LLaMA-2}_{\text{7B}}$ checkpoint from Hugging Face [Meta AI, 2023]. To examine the generalizability to attention variants and tasks, we further employ $\text{ViT}_{\text{Large}}$ [Dosovitskiy et al., 2020] and $\text{LLaVA-1.5}_{\text{7B}}$ [Liu et al., 2023].
Datasets. We evaluate on various widely-adopted benchmarks: GLUE [Wang et al., 2018], HellaSwag [Zellers et al., 2019], Winogrande [Sakaguchi et al., 2020], the AI2 Reasoning Challenge—ARC-e/ARC-c [Clark et al., 2018], OBQA [Mihaylov et al., 2018], ImageNet1k [Deng et al., 2009], CIFAR-100 [Krizhevsky, 2009], Food-101 [Bossard et al., 2014], Fashion MNIST [Xiao et al., 2017], VizWiz-VQA [Gurari et al., 2018], and MM-Vet [Yu et al., 2024].
@@
5.2 Main Results
We evaluate HIES using two key metrics: model quality and stability. [^4] Experimental details are provided in Appendix 5.1. As reported in Table 1, HIES improves model quality by 8.21% on average. Table 2 further demonstrates a 3.3% average stability gain over HIS. Notably, at a pruning ratio of 50%, HIES achieves gains of up to 15.2% in model quality and $2.04\times$ in stability compared to the best-performing baseline. These results corroborate our theoretical analysis, demonstrating that HIES preserves end-task performance while markedly enhancing robustness, both critical for reliable and efficient deployment.
@@ 5.2 Main Results
10% HIS 97.71 96.55 94.36 97.30 99.47 95.79 95.66 94.22 96.63 0.00%
L2 95.18 76.41 26.72 68.63 21.87 60.33 72.44 39.71 57.41 -40.58%
AD 98.51 96.36 93.63 97.80 89.07 97.25 96.15 94.95 95.71 -0.95%
LLM-Pruner 97.02 85.81 75.25 85.40 39.53 90.39 84.98 86.28 80.71 -16.50%
HIES (ours) 98.17 96.36 94.12 97.80 96.67 97.25 95.74 93.50 96.45 -0.19%
30% HIS 94.38 90.03 90.03 94.77 78.93 92.95 91.56 87.36 90.25 0.00%
L2 90.94 83.13 26.72 63.90 20.73 50.39 59.23 42.24 54.66 -39.45%
AD 88.53 81.21 78.68 87.90 24.53 63.04 79.07 69.68 71.58 -20.70%
LLM-Pruner 92.66 63.09 84.80 82.37 62.13 68.70 68.29 70.04 74.76 -17.21%
HIES (ours) 97.13 93.38 89.46 94.97 86.67 93.23 91.56 88.81 91.90 +1.83%
@@
L2 86.47 82.26 26.72 63.90 20.40 49.81 45.58 34.30 51.93 -32.76%
AD 52.18 81.21 51.72 72.33 21.87 73.49 68.49 62.82 60.76 -21.32%
LLM-Pruner 89.68 41.32 78.19 83.27 44.13 61.58 67.52 75.81 67.19 -13.02%
HIES (ours) 95.07 83.13 84.56 88.27 75.20 85.54 78.29 81.23 83.66 +8.34%

### v2

@@ 1 Introduction
However, HIS-based methods often exhibit limited stability in their performance. For clarity, prior works (Bair et al., 2023; Blanchet et al., 2024) motivate treating stability as a practical surrogate for robustness—namely, a model’s resilience to input perturbations and pruning-induced distributional shifts. Such stability is crucial in real-world deployments where distribution shifts are common and aggressive compression is often required. In our observations, this instability appears to stem from two key factors. First, existing HIS-based methods solely rely on the loss gradient with respect to each head’s output, which fails to capture token-level attention allocation or its alignment with the task’s empirical distribution. Consequently, a concentrated head and a diffuse head can yield similar important scores, concealing their functionally distinct roles on the task-specific data manifold. Second, a uniform, layer-agnostic criterion precludes layer-specific adaptation despite evidence that different layers require distinct attention behaviors (Artzy and Schwartz, 2024). Lacking such layer-specific characteristics often results in imbalanced pruning—preserving redundant heads in some layers while removing functionally critical ones in others. This imbalance not only degrades accuracy but also undermines stability, leading to unpredictable performance fluctuations across inputs or compression levels, particularly under aggressive pruning ratios.
@@ 2 Background
Attention head pruning. To compress large language models efficiently, structured pruning methods (Han et al., 2015; Wang et al., 2019; Xia et al., 2022; Ma et al., 2023; Ashkboos et al., 2024), which remove specific architectural components from Transformer models, have been widely adopted. Among these, attention head pruning has gained traction. This is largely because it directly reduces attention FLOPs and KV-cache memory while preserving the layer topology, thereby simplifying checkpoint compatibility and serving integration. Consequently, large‐scale studies adopt head-level pruning as a practical axis in LLM compression pipelines (Jaradat et al., 2024; Muralidharan et al., 2024) [^1] For more detailed discussions on related work, please refer to Appendix A.. Attention head pruning removes selected heads from a trained Transformer’s multi‑head attention with minimal impact on end‑task performance (Vaswani et al., 2017). A widely adopted criterion is the HIS of Michel et al. (2019), which introduces mask variables $m_{h}\in\{0,1\}$ multiplying the output of head $h$ and defines importance as the expected first-order loss increase under masking:
@@ 3 Motivation
Pruning Transformer models is most commonly driven by gradient-based criteria, such as HIS and variants used in recent pruning frameworks (e.g., LLM-Pruner) [Michel et al., 2019, Ma et al., 2023]. While gradient-based methods are often effective at moderate sparsity, they exhibit sharp accuracy degradation once the pruning ratio exceeds a certain threshold, as shown in Figure 1 (a). Such sharp-drops have been widely observed across various attention variants and tasks [Ma et al., 2023, Li et al., 2023a, Ghattas et al., 2025], underscoring the generality of this phenomenon.
We focus on this “sharp-drop” regime and contrast two groups of heads. The first group consists of low-HIS heads that are pruned during the sharp-drop of accuracy. The second group consists of high-HIS heads that remain unpruned. The attention score heatmap in Figure 1 (b) reveals that some pruned heads (red table) assign high attention scores—i.e., the weights computed by the softmax over token-token similarity that indicate how strongly a token attends to another—to sentiment-discriminative tokens (tokens relevant for classification). Nonetheless, these heads are pruned due to their low HIS values and end up causing the sharp accuracy drop observed in Figure 1 (a). In contrast, some unpruned heads (gray table) often allocate strong attention to non-informative tokens. These heads, however, have high HIS values and thus remain unpruned, though they contribute little to overall model quality. These analysis results demonstrate that the gradient-based $\mathrm{HIS}$ is insufficient to capture the token-level attention score distributions (the detailed mathematical analysis is provided in Section 4.2.1), thereby resulting in suboptimal pruning decisions for heads focusing on decisive tokens.
The sharp-drop observed in pruning can be interpreted as a collapse of structural diversity in attention behaviors, caused by the elimination of heads that concentrate on decisive tokens. To capture and prevent such collapse of structural diversity in attention, we employ attention entropy, a measure widely used to prevent policy collapse in reinforcement learning [Bharadhwaj et al., 2020, Xiao et al., 2021, Wang et al., 2025]. As shown in the bottom-right plots of Figure 1 (b), incorporating AE helps retain low-entropy heads by recognizing their concentrated focus on decisive tokens, thereby preventing them from being pruned and mitigating the sharp-drop in accuracy.
@@ 4 Proposed Method
Figure 2 provides an overview of our proposed Head Importance–Entropy Score (HIES). Figure 2 (a) outlines the Transformer architecture, where HIS computes the importance score of each of the $h$ attention heads within a Transformer block. Figure 2 (b) displays the heatmap of the first-order, loss-based HIS. Combining the normalized HIS and AE yields the HIES heatmap in Figure 2 (c), which integrates complementary signals provide a more stable assessment of attention heads across layers. This design captures the key intuition of HIES: HIS can be reinforced by AE, providing a more robust basis for pruning head selections. Section 4.1 formalizes HIES, and Section 4.2 develops a risk-decomposition analysis that clarifies the complementary roles of HIS and AE and motivates HIES as a robust criterion for head selection in pruning.
@@ 4.1 Head Importance Entropy Score
Min-Max Normalization Directly comparing raw HIS and AE is inherently problematic, as the two metrics reside on different scales and encode distinct types of signals. To enable meaningful integration and ranking, we apply min–max normalization to both metrics, rescaling their values to the interval $[0,1]$: $\widehat{\text{HIS}}_{h}=\frac{\text{HIS}_{h}-\min(\text{HIS})}{\max(\text{HIS})-\min(\text{HIS})},\quad\widehat{\text{AE}}_{h}=\frac{\text{AE}_{h}-\min(\text{AE})}{\max(\text{AE})-\min(\text{AE})}.$ This distribution-agnostic normalization improves cross-criterion interpretability; lower normalized scores denote higher pruning priority. Prior studies show min–max scaling outperforms $z$-score standardization in stability and reproducibility across diverse tasks [De Amorim et al., 2023, Lima and Souza, 2023].
@@ 4.2.1 Loss-Increase Control via Head Importance (HIS)
$\begin{gathered}\mathrm{HIS}_{h}~:=~\mathbb{E}_{x\sim\mathcal{D}}\!\left|\frac{\partial\mathcal{L}(x)}{\partial m_{h}}\right|\\ ~=~\mathbb{E}_{x\sim\mathcal{D}}\!\left|\,\big\langle\nabla_{A_{h}(x)}\mathcal{L}(x),\,A_{h}(x)\big\rangle_{F}\,\right|\\ ~=~\mathbb{E}_{\mathcal{D}}\!\Bigl[\,\left|\,\langle\nabla_{A_{h}}\mathcal{L},\;A_{h}\rangle_{F}\,\right|\,\Bigr].\end{gathered}$ (1)
@@
Lemma 4.1 (Loss-increase upper bound under head masking).
Let $\beta_{y}:=\|\nabla_{\mathbf{y}}^{2}\mathcal{L}\|_{2}$. For any mask variables $\{m_{h}\}_{h=1}^{|\mathrm{H}|}$,
$\Delta\mathcal{L}\coloneqq\mathbb{E}_{\mathcal{D}}\!\bigl[\mathcal{L}(\mathbf{y}+\delta\mathbf{y})-\mathcal{L}(\mathbf{y})\bigr]\;\leq$ (2)
$\sum_{h=1}^{|\mathrm{H}|}(1-m_{h})\,\mathrm{HIS}_{h}+\frac{\beta_{y}}{2}\sum_{h=1}^{|\mathrm{H}|}(1-m_{h})\,\|A_{h}\|_{F}^{2}.$
Moreover, under binary (sigmoid) cross-entropy we have $\|\nabla_{\mathbf{z}}^{2}\mathcal{L}\|_{2}\leq\tfrac{1}{4}$; with the linear projection $\mathbf{z}=\mathbf{y}W^{O}$, this yields $\beta_{y}\leq\tfrac{1}{4}\,\|W^{O}\|_{2}^{2}$, hence
@@
Eq. (3) shows that, for a fixed pruning fraction $\rho=\tfrac{1}{\mathrm{|H|}}\sum_{h}(1-m_{h})$, selecting heads with the smallest $\mathrm{HIS}_{h}$ minimizes the dominant first-order term, while the quadratic term is controlled by $\|W^{O}\|_{2}$ (or blockwise norms) and token-averaged activations. Under standard normalization, the quadratic contribution is typically dominated by the first-order term (cf. Appendix B.1.5), justifying the use of $\mathrm{HIS}_{h}$ as a practical surrogate importance for head pruning.
However, since $A_{h}=\mathrm{Attn}_{h}V_{h}$, $\mathrm{HIS}_{h}$ in Eq. (1) depends solely on $A_{h}$ and $\nabla_{A_{h}}\mathcal{L}$, without explicitly incorporating the distribution $\mathrm{Attn}_{h}$. Consequently, two heads with very different attention patterns (e.g., sharply focusing on tokens versus spreading over tokens) can yield similar $\mathrm{HIS}_{h}$ whenever the resulting $A_{h}$ is comparable. This is supported by our empirical analysis in Section 3.
@@ 4.2.2 Generalization Gap and Attention Entropy (AE)
$\begin{gathered}\mathrm{AD}_{h}(x):=\frac{1}{n(x)\,\log n(x)}\sum_{t=1}^{n(x)}\Bigl(\log n(x)-H\!\bigl(\bm{\alpha}^{(h)}_{t}(x)\bigr)\Bigr)\\ \;=\;1-\mathrm{AE}_{h}(x)\in[0,1].\end{gathered}$ (4)
@@ 4.2.3 Risk Upper Bound and HIES Minimization
$\mathcal{R}(m)\;:=\;\sum_{h=1}^{|\text{H}|}(1-m_{h})\,\text{HIES}_{h}\vskip-7.11317pt$ (6)
@@
$\min_{m\in\{0,1\}^{|\text{H}|}}\;\sum_{h=1}^{|\text{H}|}(1-m_{h})\,\text{HIES}_{h}\quad\text{s.t.}\;\sum_{h=1}^{|\text{H}|}m_{h}\;=\;k.$ (7)
Lemma 4.2 (Optimality).
Selecting the $k$ heads with the largest HIES values (equivalently, pruning the $|\text{H}|-k$ heads with the smallest HIES values) yields the globally optimal mask $m^{*}$ that minimizes (6) subject to (7).
@@ 4.2.4 Orthogonality and Complementarity
$\begin{gathered}u_{h}\;:=\;\operatorname{sign}\!\bigl(\bm{\alpha}^{(h)\!\top}g_{h}\bigr)\,g_{h},v_{h}\;:=\;\mathbf{1}+\log\bm{\alpha}^{(h)},\\ \tilde{u}_{h}:=P\,u_{h},\;\;\tilde{v}_{h}:=P\,v_{h},\end{gathered}$
where $P:=I-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\!\top}$ projects onto $\{w:\mathbf{1}^{\!\top}w=0\}$. Assume $\mathrm{Cov}(\tilde{u}_{h},\tilde{v}_{h})=0$ (the cross-covariance matrix is zero). If, in addition, either $\mathbb{E}_{\mathcal{S}}[\tilde{u}_{h}]=0$ or more generally $\langle\mathbb{E}_{\mathcal{S}}[\tilde{u}_{h}],\,\mathbb{E}_{\mathcal{S}}[\tilde{v}_{h}]\rangle=0$ [^3] We provide an empirical sanity check supporting the assumption, which is consistently validated on both $\text{BERT}_{\text{base}}$ and $\text{LLaMA-2}_{\text{7B}}$ in Appendix D.1. , then
@@
i.e., the two gradient directions are orthogonal in expectation [^4] Detailed derivations and preliminaries are deferred to Appendix B.4..
@@ 5.1 Experimental Setup
Model. We use publicly available $\text{BERT}_{\text{base}}$ checkpoints that have been fine-tuned and released by prior work [Devlin et al., 2019], and $\text{LLaMA-2}_{\text{7B}}$ checkpoint from Hugging Face [Touvron et al., 2023]. To examine the generalizability to attention variants and tasks, we further employ $\text{ViT}_{\text{Large}}$ [Dosovitskiy, 2020] and $\text{LLaVA-1.5}_{\text{7B}}$ [Liu et al., 2023].
Datasets. We evaluate on various widely-adopted benchmarks: GLUE [Wang et al., 2018], HellaSwag [Zellers et al., 2019], Winogrande [Sakaguchi et al., 2021], the AI2 Reasoning Challenge—ARC-e/ARC-c [Clark et al., 2018], OBQA [Mihaylov et al., 2018], ImageNet1k [Deng et al., 2009], CIFAR-100 [Krizhevsky et al., 2009], Food-101 [Bossard et al., 2014], Fashion MNIST [Xiao et al., 2017], MMLU [Hendrycks et al., 2020], GSM8K [Cobbe et al., 2021], VizWiz-VQA [Gurari et al., 2018], and MM-Vet [Yu et al., 2023].
@@
10% HIS 97.71 96.55 94.36 97.30 99.47 95.79 95.66 94.22 96.38 0.00%
L2 95.18 76.41 26.72 68.63 21.87 60.33 72.44 39.71 57.66 -40.17%
AD 98.51 96.36 93.63 97.80 89.07 97.25 96.15 94.95 95.47 -0.94%
LLM-Pruner 97.02 85.81 75.25 85.40 39.53 90.39 84.98 86.28 80.58 -16.39%
HIES (ours) 98.17 96.36 94.12 97.80 96.67 97.25 95.74 93.50 96.20 -0.19%
30% HIS 94.38 90.03 90.03 94.77 78.93 92.95 91.56 87.36 90.00 0.00%
L2 90.94 83.13 26.72 63.90 20.73 50.39 59.23 42.24 54.66 -39.27%
AD 88.53 81.21 78.68 87.90 24.53 63.04 79.07 69.68 71.58 -20.47%
LLM-Pruner 92.66 63.09 84.80 82.37 62.13 68.70 68.29 70.04 74.01 -17.77%
HIES (ours) 97.13 93.38 89.46 94.97 86.67 93.23 91.56 88.81 91.90 +2.11%
@@
L2 86.47 82.26 26.72 63.90 20.40 49.81 45.58 34.30 51.18 -33.73%
AD 52.18 81.21 51.72 72.33 21.87 73.49 68.49 62.82 60.51 -21.65%
LLM-Pruner 89.68 41.32 78.19 83.27 44.13 61.58 67.52 75.81 67.69 -12.35%
HIES (ours) 95.07 83.13 84.56 88.27 75.20 85.54 78.29 81.23 83.91 +8.65%
5.2 Main Results
We evaluate HIES using two key metrics: model quality and stability [^5] Experimental details are provided in Appendix C. We also conduct sensitivity analysis across multiple random seeds and observe consistent improvements for both $\text{BERT}_{\text{base}}$ and $\text{LLaMA-2}_{\text{7B}}$. Details are provided in Appendix D.2. Note that we use a calibration size of 32 for computing HIES; additional analyses are reported in Appendix D.3.. As reported in Table 1, HIES improves model quality by 8.21% on average. Table 2 further demonstrates a 3.3% average stability gain over HIS. Notably, at a pruning ratio of 50%, HIES achieves gains of up to 15.2% in model quality and $2.04\times$ in stability compared to the best-performing baseline. These results corroborate our theoretical analysis, demonstrating that HIES preserves end-task performance while markedly enhancing robustness, both critical for reliable and efficient deployment.